Let tan A = 7 and tan B = 9. Which is cot A + cot B?

A. 0.06
B. 0.6
C. .126984
D. 0.25
1
Expert's answer
2013-07-11T08:22:12-0400
cot A + cot B=1/ tan A + 1/ tan B=1/7+1/9=(7+9)/63=16/63=approximately 0.25
Thus, the correct answer is D.